OpenStudy (anonymous):

how to differentiate f(x)= ln(5)x-2lnx ?

OpenStudy (reemii):

\((\ln(5))x - 2\ln x\) ?

OpenStudy (reemii):

is that the function? need to be sure about parentheses before computing. if that's the function, it will be \(f'(x) = (\ln 5) \times (x)' - 2 \times (\ln x)' = \dots\) ?